# 0.5pt (eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version

### v2

@@ Abstract
Neural Radiance Fields (NeRFs) have shown remarkable novel view synthesis capabilities even in large-scale, unbounded scenes, albeit requiring hundreds of views or introducing artifacts in sparser settings. Their optimization suffers from shape-radiance ambiguities wherever only a small visual overlap is available. This leads to erroneous scene geometry and artifacts. In this paper, we propose Re-Nerfing, a simple and general multi-stage data augmentation approach that leverages NeRF’s own view synthesis ability to address these limitations. With Re-Nerfing, we enhance the geometric consistency of novel views as follows: First, we train a NeRF with the available views. Then, we use the optimized NeRF to synthesize pseudo-views around the original ones with a view selection strategy to improve coverage and preserve view quality. Finally, we train a second NeRF with both the original images and the pseudo views masking out uncertain regions. Extensive experiments applying Re-Nerfing on various pipelines on the mip-NeRF 360 dataset, including Gaussian Splatting, provide valuable insights into the improvements achievable without external data or supervision, on denser and sparser input scenarios. Project page: https://renerfing.github.io.
Keywords:
Novel view synthesis Neural rendering Data augmentation
@@ 1 Introduction
Neural Radiance Fields (NeRFs) have revolutionized 3D scene representation and rendering, enabling unprecedented quality in synthesizing novel views from a set of images [33]. NeRF optimizes a continuous volumetric scene function in the weights of two multi-layer perceptrons (MLPs), which, when queried with rays sampled from posed cameras, generate corresponding RGB values and volume densities [33]. These principles have been applied across various tasks, simplifying content creation, rendering, and reconstruction workflows.
@@
Despite its remarkable success, NeRFs [33] have several limitations, such as slow training [34], failures when noisy or no camera poses are available [30, 5], and computationally intense rendering mostly incompatible with mobile applications [12]. While many works have been proposed to address these issues, collecting enough high-quality images with sufficient overlap to ensure successful model convergence remains a challenging task. Towards this end, researchers have explored the application of NeRFs in sparse view settings, with only a handful of images available [51, 55].
Figure 1: Examples of synthesized views by 3D Gaussian Splatting [26], PyNeRF [52], Instant-NGP [34] with and without the proposed Re-Nerfing. Our approach improves the image quality thanks to the enforcement of geometric constraints on synthesized pseudo views. Applied to the baselines, the proposed method enables the reconstruction of finer details and improves upon the shape-radiance ambiguity. Image crops show test views from the mip-NeRF 360 [3] dataset. The models were trained on the sparser setting (30 views).
@@
Moreover, complex geometries and large-scale environments often lead to artifacts due to the inability to consistently encode structures [43, 2, 15]. This is particularly severe in sparse settings, where the limited views lead to hallucinations manifesting, e.g., as floaters in the free space due to the ambiguity of shape and radiance [51, 15, 31]. To mitigate these issues, depth and structural priors have been inserted into the optimization to leverage additional geometric information [15, 43, 51]. However, reliable dense depth estimates are also hard to obtain, and using wrongly estimated depth can further hurt performance [15].
@@
In this work, we address these open issues with a simple and effective solution exploiting the inherent novel view synthesis capability of NeRF to improve its own performance. We achieve this with Re-Nerfing, a multi-stage, general pipeline that introduces structural constraints from synthetic views generated by a previously trained NeRF on the same scene. As shown in Figure 1, thanks to its ability to improve the view coverage while preserving quality discarding uncertain regions, Re-Nerfing enables improvements in both sparse and dense settings. The main contributions of this paper can be summarized as follows:
@@
  • With Re-Nerfing, we show how a NeRF’s and Gaussian Splatting’s novel view synthesis is beneficial to enhance their own rendering quality.
@@
  • We propose a simple and effective iterative augmentation technique to enhance any NeRF’s outputs by training a new NeRF with the addition of synthesized views selected to improve the scene coverage and masked to discard uncertain areas.
  • We provide insights on how the views should be selected and show how masking their uncertain regions can help mitigating a sub-optimal selection.
@@
  • We show the wide applicability of Re-Nerfing upon various novel view synthesis pipelines [26, 52, 34] improving in both denser and sparser settings.
@@ 2 Related Work
Given a dense set of images, traditional interpolation techniques based on light field sampling can synthesize photorealistic novel views [29]. In relatively sparser settings, neural networks have been used for view synthesis by exploiting the correlation of the visual appearance of an instance across different views [57, 17]. More recently, 3DGS [26] has shown great promise in view synthesis tasks by building on a previously reconstructed geometry and optimizing the scene as 3D Gaussians with spherical harmonics.
@@
NeRF Mildenhall et al. [33] introduced NeRF by combining neural implicit representations with differentiable rendering, encoding both volumetric properties and view-dependent appearance and leading to high-quality novel views. NeRF has catalyzed a myriad of works tackling various aspects of its limitations and further improving novel view synthesis [54]. Among these, Instant-NGP [34], FastNeRF [20] and others  [8, 47, 19] enable significantly faster training speeds, BARF [30], SPARF [51] and Nope-NeRF [5] cope with imperfect or absent camera poses, mip-NeRF 360 [3] deals with unbounded scenes and Nerfies [38] among others  [7, 39, 18] deals with dynamic scenes. Additionally, the work of Rössle et al. [43] delivers more consistent geometry, CoNeRF [25] enables altering the output, Panoptic Neural Fields [27] incorporates scene understanding information, Block-NeRF [48] renders city-scale scenes, and CamP [37] optimizes the camera parameters. Zip-NeRF [4] combines mip-NeRF 360 and grid-based models such as Instant-NGP to reach state-of-the-art results with fast training speeds.
@@
Sparse settings Nevertheless, highly sparse settings are particularly challenging as the standard multi-view constraints are not sufficient [51, 13]. Strong geometric information is required to reconstruct the scene and enable satisfactory novel views. Towards this end, researchers have explored the support of additional data and external models to provide geometric supervision [51, 43] or semantic knowledge [55, 41]. Such techniques enable novel view synthesis even from a single image. SPARF [51], for example, relies on a pixel correspondence network to enforce a geometrically accurate solution. PixelNeRF [55] learns a prior from multiple scenes and conditions the novel views on the few inputs available. However, by depending on external models and large amounts of training data, they eliminate the advantage of NeRF of not requiring extensive training data but just that for the instance or scene of interest.
@@
Geometric constraints Enforcing geometric consistency plays a crucial role in NeRF as it directly impacts the quality of the generated views [43]. DS-NeRF [15] supervises NeRF’s optimization with sparse depth obtained from COLMAP [44], thereby reducing overfitting and accelerating training. Rössle et al. [43] went a step further by feeding the sparse COLMAP point cloud to a depth completion network and constraining NeRF’s optimization according to the estimated depth and its associated uncertainty. Urban Radiance Fields [42] focuses on outdoor scenes where LiDAR data is available, which is used for losses that benefit surface estimation. SPARF [51] exploits multi-view geometry constraints and pixel correspondences, minimizing the re-projection error using the depth rendered by the NeRF and the pose estimates, which are optimized jointly.
@@
Figure 2: Re-Nerfing is a multi-stage framework. Compatible with any NeRF and 3DGS pipeline, it operates by first training a model with the available views (green, 1st). This model is then used to generate novel views from camera poses to improve the scene coverage (orange). Then, we compute the NeRF’s model uncertainty on such novel views (blue) and discard the uncertain regions (orange-blue). Finally, we use these masked views and the original ones to train a second model (orange, 2nd). The process can be repeated iteratively by generating the novel views with the second model (3rd stage).
@@
Data augmentation for NeRF Data augmentation aims to increase the pool of training data with samples that are variations of the existing ones (e.g., image flip), ultimately regularizing the model. As for other computer vision tasks and settings [35, 22, 28], data augmentation techniques have also been proposed for novel view synthesis. In this domain, PANeRF [1], GeoAug [9], and VM-NeRF [6] generate new views by warping existing ones via homography [1] or exploiting the NeRF’s depth estimates [9, 6]. Instead, Aug-NeRF [11] trains more robust NeRFs with worst-case perturbations of the input coordinates, the intermediate features, and the pre-rendering outputs.
@@
In this work, we enhance novel view synthesis in both denser and sparser settings by exploiting NeRF’s inherent view synthesis capabilities. Unlike existing approaches [51, 43, 55, 42], we do so in a data augmentation manner by using only the available views without extra data or additional models. We achieve this by training a baseline NeRF on the available data and using it to generate novel views that we add to the training data of a subsequent NeRF model. Specifically, we sample such views to improve the scene coverage while preserving the rays’ quality by masking out uncertain regions. In contrast to Tosi et al. [50] and Feldmann et al. [16], who trained a depth estimator with NeRF data, we train a NeRF with other NeRF data. Then, unlike other augmentation methods [1, 6, 9], our newly introduced views are fully synthesized from a NeRF model. As for standard (e.g., flipping and cropping) and more sophisticated augmentations [22, 35, 28], which can be mixed to further improve the models [28], ours can be combined with other augmentation techniques and is not an alternative to them.
3 Preliminaries
Neural Radiance Fields represent a continuous volumetric function in the weights of a neural representation, typically a Multi-Layer-Perceptron (MLP) [33], hash-grid [34] or (tri-)plane representation [18]. The spatial coordinates of rays originating from a set of cameras with known intrinsic and extrinsic parameters are mapped to their corresponding density and color values and then rendered with traditional volume rendering techniques.
The color $C(\mathbf{r})$ of a ray $\mathbf{r}=\mathbf{o}+\mathbf{d}t$ with origin $\mathbf{o}$ and direction $\mathbf{d}$ is then aggregated as the weighted sum of each color sample between the near and far bounds $[t_{n},t_{f}]$ of the ray as
$C(\mathbf{r})=\int_{t_{n}}^{t_{f}}w(t)\cdot c(t)\mathbf{d}t,$ (1)
where $w(t)$ are the volumetric rendering weights of each sample $t$ and $c(t)$ the corresponding color. The weight-terms $w(t)$ can be derived from the volumetric density $\sigma(t)$ with:
[…]
Using these rendering weights the expected termination depth $\hat{D}$ of the ray $\mathbf{r}$ can then be calculated by taking the weighted sum of each sampling location $t$ along the ray between $[t_{n},t_{f}]$ such that:
$\hat{D}=\int_{t_{n}}^{t_{f}}w(t)\cdot t\ \mathbf{d}t$ (3)
The continuous integrals in equations (1), (2), (3) are discretized following [33].
@@ 4 Method
As shown in Figure 2, the proposed Re-Nerfing operates in a multi-stage fashion. First, a baseline NeRF is trained with the available views (Section 3), then it is used to generate novel views to improve the scene coverage (Section 4.1), and lastly, a new NeRF is trained on the original and the synthetic views (Section 4.2), discarding uncertain regions of the rendered views to improve the signal’s quality. Re-Nerfing is a general framework that is compatible with any pipeline for novel view synthesis. Furthermore, Re-Nerfing follows an iterative process, so further rounds can be executed using the NeRF trained at the previous iteration as a baseline for the next one.
@@ 4.1 Augmenting via Synthesized Views
The proposed method builds on the simple, well-known observation that including more views in the training data has a positive impact on the output quality [51, 15]. In general, extra views improve the representation of details and textures while reducing ambiguities (e.g., color-radiance) and artifacts. Additional viewpoints bring value due to their contribution of new data from novel perspectives that can help disambiguate the geometric and visual properties of the scene. However, such extra views may not exist. This is not only the case in sparser settings [51], e.g., due to crowd-sourced data, but also in scenarios where all images available have already been used, and it is not possible to capture more [4].
With Re-Nerfing, we take this observation a step further and mitigate the lack of extra views by adding synthesized views to the training data. After training a baseline model $\mathbf{B}$, such as the one described in Section 3, we use it to generate novel views. While the synthesis procedure itself is standard [33], we do not just generate views for rendering or evaluation purposes as commonly done but leverage them as an intermediate data augmentation step. Specifically, we exploit the model’s view synthesis purpose to enhance the output quality of another model $\mathbf{D}$ trained at a later stage by adding more views of the same scene (Section 4.2).
Nevertheless, such synthesized views are not as good as real images, otherwise, the novel view synthesis problem would be already solved. To exploit the advantage of having more viewpoints, we not only synthesize them in the vicinity of the original views, where the displacement is small and the quality is better, but also relatively further, where the quality degrades. We illustrate such trade-offs in Figure 3. Additionally, we mitigate the quality degradation by masking out uncertain regions of the synthesized views (Section 4.2).
View coverage We seek to improve the view coverage of the captured scene or object through our augmentations. This is particularly relevant in sparser settings. Toward this end, we generate multiple novel views from camera poses interpolated across existing neighboring views. We define the hyperparameters $N$ and $\Omega$, being the augmentation factor and the set of $N-1$ interpolation factors $\alpha_{j}$, respectively. Specifically, for the simple case of equally spaced novel views:
$\Omega=\left\{\alpha_{j}\mid\alpha_{j}=\frac{j}{N},\;j=1,2,\ldots,N-1\right\}$ (4)
However, as depicted in Figure 3, equally spaced synthesized views are not ideal as the quality would degrade around $\alpha=0.5$. This is particularly relevant in sparser settings where the larger camera displacement between existing poses would lead to worse synthetic views.
@@
$\bar{\mathbf{p}}_{i,j}=(1-\alpha_{j})\mathbf{p}_{i}+\alpha_{j}\mathbf{p}_{i+1}$ (5)
is the interpolation of the position component given an interpolation factor $\alpha_{j}$. For the rotation, we compute the interpolated $\bar{\mathbf{q}}_{i,j}$ from the original pose quaternions $\mathbf{q}_{i}$ and $\mathbf{q}_{i+1}$ using SLERP [45] as follows:
$\bar{\mathbf{q}}_{i,j}=\frac{\sin((1-\alpha_{j})\theta)}{\sin(\theta)}\mathbf{q}_{i}+\frac{\sin(\alpha_{j}\theta)}{\sin(\theta)}\mathbf{q}_{i+1}$ (6)
@@
4.2 Re-Nerfing: Re-Training NeRF
After training the first model $\mathbf{B}$ and using it to generate novel views (Section 4.1), we train another NeRF $\mathbf{D}$ from scratch with the addition of the synthesized views. To improve the signal’s quality, we mask out uncertain regions of the generated views with Bayes’ Rays [23].
Uncertainty masks We seek to add a high-quality training signal to the new model $\mathbf{D}$ by means of the synthesized views. Since the quality of such views is not on par with the originals, they may display artifacts that can impact the renderings. Therefore, we want to remove such artifacts and improve the training signal. We do so by estimating the model uncertainty for each synthesized view with Bayes’ Rays [23] and masking out the uncertain regions. The method involves simulating spatially parametrized perturbations of the radiance field and using a Bayesian Laplace approximation to create a volumetric uncertainty field. This field can be rendered similarly to an additional color channel. Specifically, for each ray $\mathbf{r}$ of a novel pose $\bar{\mathbf{P}}$, we synthesize the color $C_{\mathbf{B}}(\mathbf{r})$ and estimate the corresponding uncertainty $U(\mathbf{r})$. Then, we filter out the rays with an uncertainty $U(\mathbf{r})$ higher than the threshold $\mu$.
To optimize the second model $\mathbf{D}$, we use the masked generated views as part of the training images and query their rays during the course of optimization. The intuition behind using these synthesized images is that they provide a more diverse yet reliable set of views, which allows the NeRF representation to better triangulate the underlying scene geometry.
This is especially important at the beginning of the training, where the density of the scene still varies significantly during optimization and where the core geometrical structure of the scene is being learned. As the optimization is close to convergence, the discrepancy in image quality between the pseudo-views and the original images increases, with the former becoming detrimental to the overall reconstruction, so we remove the synthesized ones from the sampling pool.

### v3

# Re-Nerfing: Improving Novel View Synthesis through Novel View Synthesis

@@ Abstract
Recent neural rendering and reconstruction techniques, such as NeRFs or Gaussian Splatting, have shown remarkable novel view synthesis capabilities but require hundreds of images of the scene from diverse viewpoints to render high-quality novel views. With fewer images available, these methods start to fail since they can no longer correctly triangulate the underlying 3D geometry and converge to a non-optimal solution. These failures can manifest as floaters or blurry renderings in sparsely observed areas of the scene. In this paper, we propose Re-Nerfing, a simple and general add-on approach that leverages novel view synthesis itself to tackle this problem. Using an already trained NVS method, we render novel views between existing ones and augment the training data to optimize a second model. This introduces additional multi-view constraints and allows the second model to converge to a better solution. With Re-Nerfing we achieve significant improvements upon multiple pipelines based on NeRF and Gaussian-Splatting in sparse view settings of the mip-NeRF 360 and LLFF datasets. Notably, Re-Nerfing does not require prior knowledge or extra supervision signals, making it a flexible and practical add-on.
@@ 1 Introduction
Novel view synthesis (NVS) methods based on Neural Radiance Field (NeRF) [37] and 3D Gaussian Splatting (3DGS)  [28] have recently revolutionized 3D scene representation and rendering, enabling unprecedented quality in synthesizing novel views from a set of images. These techniques have been applied across various tasks, simplifying content creation, rendering, and reconstruction workflows.
Figure 1: Examples of synthesized views by 3DGS [28], Instant-NGP [38], RegNeRF  [40], with and without the proposed Re-Nerfing. Our approach improves NVS by improving the optimization, reaching better global solutions thanks to the addition of novel views synthesized by the baseline model to the training data. Image crops show test views from the mip-NeRF 360 [3] dataset for 3DGS, iNGP trained on 30 views and from the LLFF [36] dataset for RegNeRF trained on 3 views.
@@
Despite its remarkable success, NeRF [37] has several limitations, such as slow training [38], failures when noisy or no camera poses are available [33, 4], and computationally intense rendering mostly incompatible with mobile applications [12]. Many works have been proposed to address these issues, such as 3DGS [28] changing representation to shorten training times and simplifying inference. However, collecting enough high-quality images with sufficient overlap to ensure successful model convergence remains a challenging task. Towards this end, researchers have explored the application of NVS in sparse view settings, with only a handful of images available [58, 40, 66].
@@
Complex geometries and large-scale environments often lead to artifacts due to the inability to correctly triangulate the scenes’ 3D structure  [49, 2, 15]. This is particularly severe in sparse settings, where the limited views lead to hallucinations, e.g. floaters in free space due to the ambiguity of shape and radiance [58, 15, 34]. To mitigate these issues, depth and structural priors have been introduced to the optimization to leverage additional geometric information [15, 49, 58]. However, reliable dense depth estimates are also hard to obtain, and using wrongly estimated depth can further hurt performance [15].
@@
In this work, we mitigate these open issues with a simple and effective add-on solution exploiting the inherent novel view synthesis capability of NVS methods. We achieve this with Re-Nerfing, a multi-stage, general pipeline that introduces structural constraints from synthetic views generated by a previously trained model on the same scene. As shown in Figure 1, thanks to its ability to improve the view coverage while preserving quality discarding uncertain regions, Re-Nerfing significantly improves the synthesis of new views. The main contributions of this paper can be summarized as follows:
@@
  • With Re-Nerfing, we show how NeRF’s and Gaussian Splatting’s novel view synthesis is beneficial to enhance their own rendering quality.
@@
  • We propose a simple and effective iterative augmentation technique to enhance any NVS outputs by training a new NVS model with the addition of synthesized views selected to improve the scene coverage and masked to discard uncertain areas.
@@
  • We show the wide applicability of Re-Nerfing upon various NVS pipelines [28, 59, 38, 40] and datasets [3, 36].
@@ 2 Related Work
Given a dense set of images, traditional interpolation techniques based on light field sampling can synthesize photorealistic novel views [31]. In relatively sparser settings, neural networks have been used for view synthesis by exploiting the correlation of the visual appearance of an instance across different views [64, 19]. More recently, 3DGS [28] has shown great promise in view synthesis tasks by optimizing the scene as 3D Gaussians with spherical harmonics.
@@
NeRF Mildenhall et al. [37] introduced NeRF by combining neural implicit representations with differentiable rendering, encoding both volumetric properties and view-dependent appearance and leading to high-quality novel views. NeRF has catalyzed a myriad of works tackling various aspects of its limitations and further improving NVS [61]. Among these, Instant-NGP [38], FastNeRF [22] and others [8, 54, 20] enable significantly faster training speeds, BARF [33], SPARF [58] and Nope-NeRF [4] cope with imperfect or absent camera poses, mip-NeRF 360 [3] deals with unbounded scenes and Nerfies [43] among others  [7, 44, 21] deals with dynamic scenes. Additionally, the work of Rössle et al. [49] delivers more consistent geometry, CoNeRF [27] enables altering the output, Panoptic Neural Fields [29] incorporates scene understanding information, Block-NeRF [55] renders city-scale scenes, and CamP [45] optimizes the camera parameters. PyNeRF [59] combines the benefits of mip-NeRF [2] and grid-based models, such as Instant-NGP, to facilitate fast training and rendering of anti-aliased images.
@@
Sparse settings Nevertheless, highly sparse settings are particularly challenging as the standard multi-view constraints are not sufficient [58, 13]. Strong geometric information is required to reconstruct the scene and enable satisfactory novel views synthesis. Towards this end, researchers have explored the support of additional data and external models to provide geometric supervision [58, 49] or semantic knowledge [62, 47]. Such techniques enable NVS even from a single image. SPARF [58], for example, relies on a pixel correspondence network to enforce a geometrically accurate solution. PixelNeRF [62] learns a prior from multiple scenes and conditions the novel views on the few inputs available. For Gaussian Splatting-based methods, multiple extensions have been proposed to deal with sparse view settings [66, 32, 41, 17]. However, all aforementioned methods depend on external models and large amounts of training data, which eliminates the advantage of NeRF and 3DGS of not requiring any training data other than that for the object or scene of interest.
@@
Geometric constraints Enforcing geometric consistency plays a crucial role in NVS as it directly impacts the quality of the generated views [49]. DS-NeRF [15] supervises NeRF’s optimization with sparse depth obtained from COLMAP [51], thereby reducing overfitting and accelerating training. Rössle et al. [49] went a step further by feeding the sparse COLMAP point cloud to a depth completion network and constraining NeRF’s optimization according to the estimated depth and its associated uncertainty. Urban Radiance Fields [48] focuses on outdoor scenes where LiDAR data is available, which is used for losses that benefit surface estimation. FSGS [66] and related methods [41, 32] on the other hand rely on monocular depth prediction to acquire a good geometry initialisation and supervision signal. SPARF [58] exploits multi-view geometry constraints and pixel correspondences, minimizing the re-projection error using the depth rendered by the NeRF and the pose estimates, which are optimized jointly.
@@
Figure 2: Re-Nerfing is a multi-stage framework. Compatible with any NVS pipeline, it operates by first training a model with the available views (green, 1st). This model is then used to generate novel views from camera poses to improve the scene coverage (orange). Then, we compute the model’s uncertainty on such novel views (blue) and discard the uncertain regions (orange-blue). Finally, we use these masked views and the original ones to train a second model (orange, 2nd). The process can be repeated iteratively by generating the novel views with the second model (3rd stage).
@@
Data augmentation for NVS In general, data augmentation aims to increase the pool of training data with samples that are variations of the existing ones (e.g., image flip), ultimately regularizing the model. As for other computer vision tasks and settings [39, 24, 30], data augmentation techniques have also been proposed for NVS. In this domain, PANeRF [1], GeoAug [9], and VM-NeRF [5] generate new views by warping existing ones via homography [1] or exploiting the NeRF’s depth estimates [9, 5]. In contrast, Aug-NeRF [11] trains more robust NeRFs with worst-case perturbations of the input coordinates, the intermediate features, and the pre-rendering outputs.
@@
In this work, we enhance NVS in sparser settings by exploiting the NVS method’s inherent view synthesis capabilities. Unlike existing approaches [58, 49, 62, 48], we do so in a data augmentation manner by using only the available views without extra data or additional models. We achieve this by training a baseline NVS method on the available data and using it to generate novel views that we add to the training data of a subsequent NVS model. Specifically, we sample such views to improve the scene coverage while preserving the images’ quality by masking out uncertain regions. In contrast to Tosi et al. [57] and Feldmann et al. [18], who trained a depth estimator with NeRF data, we train an NVS model with other NVS data. Then, unlike other augmentation methods [1, 5, 9], our newly introduced views are fully synthesized from an NVS model.
@@ 3 Method
As shown in Figure 2, our proposed Re-Nerfing operates in a multi-stage fashion. First, a baseline method is trained with the available views (Section 3.1), then it is used to generate novel views to improve the scene coverage (Section 3.2), and lastly, a new model is trained on the original and the synthetic views (Section 3.3), discarding uncertain regions of the rendered views to improve the signal’s quality. Re-Nerfing is a general framework that is compatible with any pipeline for NVS. Furthermore, Re-Nerfing follows an iterative process, so further rounds can be executed using the model trained at the previous iteration as a baseline for the next one.
3.1 NeRF and 3DGS Optimization
To better understand the shortcomings of NeRF [37] and 3DGS [28] optimizations in a sparser setting, we review how they optimize their scene representation and render images from it. Both methods are optimized through photometric consistency, where the color of each ray or pixel is compared to the ground truth color with an L2 loss or a mixture of L1 and SSIM losses [28]. This color is obtained through volumetric rendering via the following equations, where density $o$, Transmittance $T$, and color $\mathbf{c}$ are taken along a ray or z-value at intervals $\delta_{i}$:
$C=\sum^{N}_{i}T_{i}\alpha_{i}\mathbf{c}_{i},\vspace{-0.2cm}$ (1)
where
$\alpha_{i}=(1-\text{exp}(-o_{i}\delta_{i}))\;\;\;\text{and}\;\;\;T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (2)
This formulation holds for both NeRF-based architectures and neural point-based rendering approaches, such as 3DGS [28], where the image formation process is the same but images are rendered through ray-tracing or rasterization.
To represent the geometry of the given scene only the density $o_{i}$ at sampling or point locations can be optimized. If very few images of the scene are given, this density can no longer be correctly triangulated since there are not enough multi-view constraints and too much ambiguity in the color supervision. This can be seen in the simplified example in the inset, where the green cameras represent the original ones and the shaded green region represents valid solutions where density could be distributed. The overall optimization process would converge to one of the many valid solutions within the green region, i.e., a local minimum, leading to dissatisfying geometry reconstruction and poor NVS.
@@ 3.2 Augmenting via Synthesized Views
As introduced above, sparse view reconstruction is a highly under-constrained problem. In scenarios without enough views, e.g., due to crowd-sourced or previously captured images, NVS is particularly difficult. Various methods tackled this with prior knowledge [58, 15] or regularization [40]. Instead, we propose to leverage the inherent reconstruction and rendering capabilities of NVS methods to render novel views and utilize these rendered images as additional training data in subsequent optimizations (Section 3.3), thereby introducing additional multi-view constraints without prior knowledge or external models.
We motivate this approach with the simplified example in the inset figure (Section 3.1, where the orange cameras represent novel views rendered after the first optimization, and the green cameras are the originally available views. By introducing these additional views, possible solutions reduce from the green-shaded region to the orange-shaded region, making triangulation of the scene’s geometry/density easier. We then re-start the optimization from scratch (Section 3.3). Thanks to the additional viewpoints present in the training data, we can avoid some of the previously present local minima and converge to a better global solution for the given NVS method.
View coverage We seek to improve the view coverage of the captured scene or object through our augmentations. This is particularly relevant in sparser settings. Toward this end, we generate multiple novel views from camera poses interpolated across existing neighboring views. We define the hyperparameters $N$ and $\Omega$, being the augmentation factor and the set of $N-1$ interpolation factors $\beta_{j}$, respectively. Specifically, for the simple case of equally spaced novel views:
$\Omega=\left\{\beta_{j}\mid\beta_{j}=\frac{j}{N},\;j=1,2,\ldots,N-1\right\}$ (3)
However, equally spaced synthesized views are not ideal as the quality would degrade around $\beta=0.5$. This is particularly relevant in sparser settings where the larger camera displacement between existing poses would lead to worse synthetic views. This is illustrated in the Supplementary Material.
@@
$\bar{\mathbf{p}}_{i,j}=(1-\beta_{j})\mathbf{p}_{i}+\beta_{j}\mathbf{p}_{i+1}$ (4)
is the interpolation of the position component given an interpolation factor $\beta_{j}$. For the rotation, we compute the interpolated $\bar{\mathbf{q}}_{i,j}$ from the original pose quaternions $\mathbf{q}_{i}$ and $\mathbf{q}_{i+1}$ using SLERP [52] as follows:
$\bar{\mathbf{q}}_{i,j}=\frac{\sin((1-\beta_{j})\theta)}{\sin(\theta)}\mathbf{q}_{i}+\frac{\sin(\beta_{j}\theta)}{\sin(\theta)}\mathbf{q}_{i+1}$ (5)
@@
Image Quality Our proposed method has to balance the synthesized image quality with achieving as wide a baseline as possible with regard to the original views. This is because novel views close to the training views will be of higher quality and exhibit less artifacts compared to ones which are further away but add less multi-view constraints. In practice we achieve this trade-off through generating many interpolated views between the original training views $\mathbf{P}_{i}$ and $\mathbf{P}_{i+1}$ and achieve a good balance between both factors. We illustrate the trade-off in the supplementary material.
3.3 Re-Nerfing: Re-Training NeRF or 3DGS
After training the first model $\mathbf{B}$ and using it to generate novel views (Section 3.2), we train another NVS method $\mathbf{D}$ from scratch with the addition of the synthesized views. To improve the signal’s quality, we mask out uncertain regions of the generated views with Bayes’ Rays [25], where possible.
Uncertainty masks We seek to add a high-quality training signal to the new model $\mathbf{D}$ by means of the synthesized views. Since the quality of such views is not on par with the originals, they may display artifacts that can impact the renderings. Therefore, we want to remove such artifacts and improve the training signal. One way of estimating the reconstruction uncertainty is to use Bayes’Rays [25], which can be used to mask out more uncertain regions from our synthesized images. Bayes’ Rays [25] outputs a volumetric uncertainty field for every input coordinate $\mathbf{x}$:
$U(\mathbf{x})=\text{Trilinear}(\mathbf{x},\sigma),\;\;\;\text{where}\;\;\;\sigma=||\bm{\sigma}||_{2}$ (6)
Here $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ represents the diagonal entries of the variance $\bm{\Sigma}$ of the deformation field, which intuitively encodes how much one can permute the underlying NeRF geometry without affecting reconstruction quality. $\bm{\Sigma}$ is approximated via the diagonal entries of the Hessian $\bm{H}(\bm{\theta})$, which is approximated for all rays $\bm{r}$ of a sampling pool $R$ with respect to the parameters $\bm{\theta}$ of the deformation field as:
$\bm{H(\theta)}\;\approx\;\frac{2}{R}\;\sum_{\bm{r}}\bm{J_{\theta}}(\bm{r})^{T}\bm{J_{\theta}}(\bm{r})+2\lambda\bm{I}$ (7)
$\bm{\Sigma}\;\approx\;\text{diag}(\bm{H}(\bm{\theta}))^{-1}$ (8)
For the full derivation of this formulation we refer the reader to the original publication of Bayes’ Rays [25].
The uncertainty obtained at each sampling location $\mathbf{x}$ can then be rendered via volumetric rendering. Specifically, for each ray $\mathbf{r}$ of a novel pose $\bar{\mathbf{P}}$ with ray samples $\mathbf{x}$, we synthesize the color $C_{\mathbf{B}}(\mathbf{r})$ and render the corresponding uncertainty $U(\mathbf{r(x)})$. Then, we create pixel-masks where rays with $U(\mathbf{r(x)})$ higher than a threshold $\mu$ are masked out. To optimize the second model $\mathbf{D}$, we use the generated views as part of the training images and query their masked rays during the course of optimization.
Training Schedule For both NeRF and 3DGS the scene geometry is reconstructed in the early optimization stages and later they are optimized for appearance. Since the synthesized views are not as good as real images, especially when generated at positions further away from the training views, they can become a sub-optimal signal when optimizing for appearance. Therefore, we sample from our synthesized images only in the early stages of the optimization, where they help disambiguating the scene geometry, and then remove them from the training as they start to become detrimental to the appearance optimization.